OpenStudy (anonymous):

Can someone help me, please? The function below shows the number of car owners f(t), in thousands, in a city in different years t: f(t) = 1.1t2 − 2.5t + 1.5 The average rate of change of f(t) from t = 3 to t = 5 is ______ thousand owners per year.

OpenStudy (anonymous):

Will medal and fan!

OpenStudy (anonymous):

5-3=2

OpenStudy (anonymous):

substitute 2 as t in the equation

OpenStudy (anonymous):

Okay give me one second

OpenStudy (anonymous):

Would it be 0.9?

OpenStudy (anonymous):

yap. is that in your option ?

OpenStudy (anonymous):

You fill it in yourself, it isn't a multiple choice question.

OpenStudy (anonymous):

well, i think its 0.9

OpenStudy (anonymous):

I'll tell you if it's right or not when I get my grade ^.^ Thank you for your help!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you possibly help me with another question?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The length of a rectangle is represented by the function L(x) = 4x. The width of that same rectangle is represented by the function W(x) = 7x2 − 4x + 2. Which of the following shows the area of the rectangle in terms of x? (L + W)(x) = 7x2 + 2 (L + W)(x) = 7x2 − 8x + 2 (L ⋅ W)(x) = 28x3 − 16x2 + 8x (L ⋅ W)(x) = 28x3 − 4x + 2

OpenStudy (anonymous):

what is the formular for area of a rectangle?

OpenStudy (anonymous):

A=w*l?

OpenStudy (anonymous):

open a new question and re ask

OpenStudy (anonymous):

Okay~

OpenStudy (anonymous):

i am waiting. don't forget to tag me